# Both Direct and Indirect Evidence Contribute to Dative Alternation Preferences in Language Models

### v2

@@ 4.1 Experiment 1a: Effect of Balancing Dative Alternating Forms
We balance the dataset by creating alternate forms for each occurrence of a dative, using the method outlined in § 2.5. For instance, for the DO sentence I gave the dog a bone we create its PO form I gave a bone to the dog and then add both sentences in the training set (see Table 1). We ensure that the LMs trained on this manipulated corpus still have roughly the same number of datives in total as default after artificial pollution by balancing 32,850 attested DOs and 32,850 attested POs with their unattested alternants. By design, this model should observe no word-order preferences in datives. We refer to this model as balanced. As can be seen in Fig. 3, the length effect is weakened in the balanced results but still very much present: the correlation goes from $r=-0.43$ in the default to $r=-0.33$. As shown in Table 4, our global regression model showed that the effect of balancing significantly weakened the length and animacy effects relative to default. But, as shown in Table 2, using the same per-model regression as before but focusing on the balanced model’s data, we again found a significant effect of length (but not for animacy). We also once again found verb-specific animacy effects.
4.2 Experiment 1b: Effect of Reversing All Datives
@@
[…]
This is indeed the case: swapped-datives has a roughly neutral length effect of $r=0.03$. Animacy effects are reversed, which shows that swapping direct evidence has an effect on dative preferences. swapped-datives showed significantly smaller length and animacy effects than the default models (see Table 4). Length and animacy coefficients were not significant by our per-condition regression analysis (see Table 2). Thus the swapped-datives models, as expected, finds effects in the opposite direction of default, but the magnitude is smaller.
@@ 4.3 Experiment 1c: Effect of Removing Dative and Ditransitive-like Constructions
As with balanced datives, the length and animacy effects are reduced compared to the default model. As shown in Fig. 3 and Table 2, we again find a persistent length effect: $r=-0.24$ for no-datives and $r=-0.22$ for no-2postverbal. While there is not significant evidence for a global animacy-first preference, we again observe verb-specific effects that are interestingly similar to the effects in the default case (see Fig. 6 in App. E).
@@ 4.4 Interim Discussion
Overall, two different methods of changing the input data (balancing and removing datives) fail to remove the length effect, although in all cases, the effect is unsurprisingly weaker than in the default model. Only when all the input datives are in reversed order do length effects become closer to neutral (though interestingly they are not reversed). The effects of animacy are more nuanced. Here, we do not find a global “animate-first” preference—instead, this preference seems to depend on the verb (see App. E). For instance, when the recipient is more animate, leave and sell are strongly preferred in the PO, while bring and write show strong DO preferences. At the same time, the verb-specific effects are similar between the default and no-datives condition (Fig. 6), suggesting some higher order effect of indirect evidence. We leave deeper exploration of this to future work.
@@ 5 Experiment 2: Effect of Global Length Manipulation
To perform our global length manipulations, we operate on our no-2postverbal corpus that excludes datives and related constructions. For each sentence, the children of any node in its dependency representation are ordered by length (either short-first or long-first). Since a long-first preference has been observed in head-final languages (Yamashita & Chang, 2001; Futrell et al., 2020), we also create a corpus where constituents are arranged by decreasing length and every head appears in the final position. See Algorithm 1 for pseudocode. We train a short-first, long-first, and long-first-headfinal LM on these corpora, respectively. As a control, we also train a model on a corpus where child-constituent balancing is randomly chosen (random-first). The training sets for these four models are identical apart from constituent ordering, and so no counterfactual pollution is used.
The training set of no-2postverbal is already strongly short-first – roughly 70% of entries are short-first in their original form and 44% long-first. Disregarding simple sentences where no head has more than one child (since these are automatically short/long-first), 53% of sentences are already short-first compared to 16% for long-first. This suggests that preferences learned by the short-first model should be similar to that of no-2postverbal. We provide a more detailed analysis of how short-first the original corpus is in App. I.
@@ 5.1 Results
As shown in Fig. 3, short-first, random-first, long-first, and long-first-headfinal show a length-effect gradient, appearing in the expected order. To assess significance of these results, we ran another regression (similar to our global regression, see App. C for details and Table 5 for results), focused on just the length-manipulated conditions and with the random-first condition as baseline. As expected, the long-first and long-first-headfinal models showed significantly weakened length preferences compared to random-first, but short-first showed significantly stronger length preferences than random-first. Therefore, when no direct evidence is available, the learned dative preferences tend towards preferences in the hypothesized indirect evidence.
Since the training data in many of our experiments are heavily manipulated, it is possible that observed length effects are due to higher perplexities from less good models overall. That is, maybe some models just don’t learn dative preferences very well because they don’t learn anything very well. To test this, we compute the geometric mean of perplexities on 10,000 sentences from the validation set with corresponding length manipulations for every model. These perplexity values and length effect coefficients per seed are shown in Fig. 4.
@@ 6 Conclusion
Bespoke small models learn reasonably human-like preferences for the dative alternation. Length preference is persistent even when we train models on datasets without direct evidence for such a preference, suggesting that a short-first preference is a more general property of language. By then training models on counterfactual languages where the indirect evidence become decreasingly “short-first”, we show that length effect is gradually lost, indicative of the role of indirect evidence in acquiring that preference. The story for animacy is more subtle. Instead of a global animate-first effect, we found verb-specific differences where other factors such as the theme might also play a role. At the same time, these verb-specific effects were persistent even in the absence of direct evidence, and oftentimes did not flip in ablations that flipped the direct evidence.
@@ Acknowledgments
We thank John Beavers, Sasha Boguraev, Richard Futrell, and the UT Computational Linguistics research community for discussions and feedback. Q.Y. thanks Simon Todd for his mentorship during his undergraduate thesis, which extended into this work. Q.Y. is supported by the Harrington Fellowship at UT Austin. L.W. was supported by a postdoctoral fellowship of the German Academic Exchange Service (DAAD).
@@ Appendix A Training Details
Our models are trained largely following the procedure in Misra & Mahowald (2024). We have a total of eleven model-types, of which nine are discussed in the main text (with two following a more strict definition of datives discussed in App. D), and for each type, we conducted three training runs with different seeds to account for variation due to initialization. In total, this amounts to 33 LM training runs. See Table 3 for a summary of training details—hyperparameters, compute hardware, etc.
@@ C.2 Global Regression Models
In Table 4, the interaction coefficients (e.g., balanced : Length Difference) can be thought of as telling us how much the Length or Animacy difference differs in the given condition, as opposed to the default baseline.
@@
Intercept $-0.459$***
balanced $0.046$***
no-datives $0.078$***
no-2postverbal $0.087$***
swapped-datives $0.121$***
short-first $0.127$***
random-first $0.139$***
long-first $0.249$***
long-first-headfinal $0.248$***
Length Difference $-0.189$***
Animacy Difference $0.091$***
balanced : Length Difference $0.105$***
no-datives : Length Difference $0.135$***
no-2postverbal : Length Difference $0.140$***
swapped-datives : Length Difference $0.227$***
short-first : Length Difference $0.131$***
random-first : Length Difference $0.187$***
long-first : Length Difference $0.225$***
long-first-headfinal : Length Difference $0.266$***
balanced : Animacy Difference $-0.117$***
no-datives : Animacy Difference $-0.104$***
no-2postverbal : Animacy Difference $-0.109$***
swapped-datives : Animacy Difference $-0.195$***
short-first : Animacy Difference $-0.070$***
random-first : Animacy Difference $-0.066$***
long-first : Animacy Difference $-0.034$*
long-first-headfinal : Animacy Difference $-0.126$
@@
long-first-headfinal $0.109$***
@@
short-first : Length difference $0.055$***
long-first : Length Difference $0.039$***
long-first-headfinal : Length Difference $0.080$***
@@
long-first : Animacy Difference $0.032$*
long-first-headfinal : Animacy Difference $0.060$***
@@ Appendix D Strict-dative default and balanced Results
[…]
@@
Both the regression coefficients in Table 6 and the length and animacy effects in Fig. 5 for the strict models are similar to their loose counterparts. We therefore focus on just the loose notion in the main paper, so that the ablated models are more aggressively removing direct evidence.
@@ Appendix E Verb-specific Random Slopes for Animacy
Random slopes by verb lemma for default, balanced, swapped-datives, and no-datives are shown in Fig. 6. We see that animacy effects for certain verbs are heavily affected by the manipulations in balanced and swapped-datives, indicating that the animacy preferences of these verbs are more tied to direct evidence. However, no-datives recovers similar, albeit a bit weaker, animacy effects as in default. This suggests that usages of dative verbs in the indirect evidence can inform an LM’s dative animacy preferences.
Figure 6: Verb-specific random slopes for the animacy term in our linear mixed-effects model analysis. Positive value indicates greater DO preference for animate recipients, while negative ones indicate low DO preference for inanimate recipients.
@@ Appendix F Detecting Datives and 2postverbals
Using spacy, we first detect all utterances containing a verb which has a dative and a dobj dependency or two dobj dependencies as possible DO datives, and verbs with a dobj dependency, and with either a dative or prep arc to to with a pobj or a dative arc to for with a pobj as possible PO datives. From these entries, we perform a sanity check to ensure that the noun phrases come after the verb, and the recipient and theme have appropriate parts of speech.
These entries are then loose-datives. From these entries, we check if the verb lemma is classified as a dative or benefactive verb in Levin (1993). If so, we check if it is in the permissible form (e.g. if the verb is classified as benefactive only, the entry cannot be a to-PO, or if the verb is classfied as DO only, the entry cannot be a PO). If a loose-dative satisfies the above, it is also a strict-dative.
@@ Appendix G Artificial Pollution
The counts of dative exposure by model are listed in Table 7.
@@ Appendix I Measuring the Short-Firstness of English
For each sentence, we compute the inversion number corresponding a head that has $n\geq 2$ dependents as the number of out-of-order pairs (w.r.t. to short-first or long-first). The inversion number of a head is precisely equal to the number of adjacent swaps to sort its dependents. As such, a head with $n$ dependents has an inversion number at most $\binom{n}{2}$.
@@
For entries where constituent rearrangements were possible (i.e. there exists one head with two or more dependents), it takes 23.8% of the maximal number of swaps to reach short-first, compared to 53.2% for long-first. This means that on average, it takes more than twice the number of adjacent-swap operations to sort by long-first compared to short-first.

### v3

@@ 4.1 Experiment 1a: Effect of Balancing Dative Alternating Forms
Table 2: Regression coefficients from mixed effect models, fit separately for each LM’s DO preference, with * for $p<.05$, ** for $p<0.01$, and *** for p<.001{p<.001}. Length and animacy are difference scores, as described in §˜2.5.
We balance the dataset by creating alternate forms for each occurrence of a dative, using the method outlined in §˜2.5. For instance, for the DO sentence I gave the dog a bone we create its PO form I gave a bone to the dog and then add both sentences in the training set (see Table˜1). We ensure that the LMs trained on this manipulated corpus still have roughly the same number of datives in total as default after artificial pollution by balancing 32,850 attested DOs and 32,850 attested POs with their unattested alternants. By design, this model should observe no word-order preferences in datives. We refer to this model as balanced. As can be seen in Fig.˜4, the length effect is weakened in the balanced results but still very much present: the correlation goes from r=0.43r=-0.43 in the default to r=0.33r=-0.33. As shown in Table˜4, our global regression model showed that the effect of balancing significantly weakened the length and animacy effects relative to default. But, as shown in Table˜2, using the same per-model regression as before but focusing on the balanced model’s data, we again found a significant effect of length (but not for animacy). We also once again found verb-specific animacy effects.
4.2 Experiment 1b: Effect of Reversing All Datives
This is indeed the case: swapped-datives has a roughly neutral length effect of $r=0.03$. Animacy effects are reversed, suggesting swapping direct evidence has an effect on dative preferences. swapped-datives showed significantly smaller length and animacy effects than the default LMs (see Table˜4). Length and animacy coefficients were not significant by our per-condition regression analysis (see Table˜2). Thus the swapped-datives models, as expected, finds effects in the opposite direction of default, but the magnitude is smaller.
@@ 4.3 Experiment 1c: Effect of Removing Dative and Ditransitive-like Constructions
As with balanced datives, the length and animacy effects are reduced compared to the default model. As shown in Fig.˜4 and Table˜2, we again find a persistent length effect: r=0.24r=-0.24 for no-datives and r=0.22r=-0.22 for no-2postverbal. While there is not significant evidence for a global animacy-first preference, we again observe verb-specific effects that are interestingly similar to the effects in the default case (see Fig.˜8 in App.˜E).
@@ 4.4 Interim Discussion
Overall, two different methods of changing the input data (balancing and removing datives) fail to remove the length effect, although in all cases, the effect is unsurprisingly weaker than in the default model. Only when all the input datives are in reversed order do length effects become closer to neutral (though interestingly they are not reversed). The effects of animacy are more nuanced. Here, we do not find a global “animate-first” preference—instead, this preference seems to depend on the verb (see App.˜E). For instance, when the recipient is more animate, leave and sell are strongly preferred in the PO, while bring and write show strong DO preferences. At the same time, the verb-specific effects are similar between the default and no-datives condition (Fig.˜8), suggesting some higher order effect of indirect evidence. We leave deeper exploration of this to future work.
@@ 5 Experiment 2: Effect of Global Length Manipulation
To perform our global length manipulations, we operate on our no-2postverbal corpus that excludes datives and related constructions. For each sentence, the children of any node in its dependency representation are ordered by length (either short-first or long-first). Since a long-first preference has been observed in head-final languages (Yamashita & Chang, 2001; Futrell et al., 2020), we also create a corpus where constituents are arranged by decreasing length and every head appears in the final position. See Algorithm˜1 for pseudocode. We train a short-first, long-first, and long-first-headfinal LM on these corpora, respectively. As a control, we also train a model on a corpus where child-constituent balancing is randomly chosen (random-first). The training sets for these four models are identical apart from constituent ordering, and so no counterfactual pollution is used.
The training set of no-2postverbal is already strongly short-first – roughly 70% of entries are short-first in their original form and 44% long-first. Disregarding simple sentences where no head has more than one child (since these are automatically short/long-first), 53% of sentences are already short-first compared to 16% for long-first. This suggests that preferences learned by the short-first model should be similar to that of no-2postverbal. We provide a more detailed analysis of how short-first the original corpus is in App.˜I.
@@
Figure 5: Correlation between DO preference and length difference (top) and animacy difference (bottom) by model on instances in the test set ($N=2000$). Models shown here are for Experiment 2.
@@ 5.1 Results
As shown in Fig.˜5, short-first, random-first, long-first, and long-first-headfinal show a length-effect gradient, appearing in the expected order. To assess significance of these results, we ran another regression (similar to our global regression, see App.˜C for details and Table˜5 for results), focused on just the length-manipulated conditions and with the random-first condition as baseline. As expected, the long-first and long-first-headfinal models showed significantly weakened length preferences compared to random-first, but short-first showed significantly stronger length preferences than random-first. Therefore, when no direct evidence is available, the learned dative preferences tend towards preferences in the hypothesized indirect evidence.
Since the training data in many of our experiments are heavily manipulated, it is possible that observed length effects are due to higher perplexities from less good models overall. That is, maybe some models just don’t learn dative preferences very well because they don’t learn anything very well. To test this, we compute the geometric mean of perplexities on 10,000 sentences from the validation set with corresponding length manipulations for every model. These perplexity values and length effect coefficients per seed are shown in Fig.˜6.
@@ 6 Conclusion
Bespoke small models learn reasonably human-like preferences for the dative alternation. Length preference is persistent even when we train models on datasets without direct evidence for such a preference, suggesting that a short-first preference is a more general property of English. By then training models on counterfactual languages where the indirect evidence becomes decreasingly “short-first”, we show that the length effect is gradually lost, indicative of the role of indirect evidence in acquiring that preference. The story for animacy is more subtle. Instead of a global animate-first effect, we found verb-specific differences where other factors such as the theme might also play a role. At the same time, these verb-specific effects were persistent even in the absence of direct evidence, and oftentimes did not flip in ablations that flipped the direct evidence.
@@ Acknowledgments
We thank John Beavers, Sasha Boguraev, Richard Futrell, and the UT Computational Linguistics research community for discussions and feedback. Qing Yao thanks Simon Todd for his mentorship during his undergraduate thesis, which extended into this work. Qing Yao is supported by the Donald D. Harrington Fellowship at UT Austin. Kanishka Misra is supported by the Donald D. Harrington Faculty Fellowship at UT Austin. Leonie Weissweiler was supported by a postdoctoral fellowship of the German Academic Exchange Service (DAAD). Kyle Mahowald was supported by NSF CAREER grant 2339729 from the Director of STEM Education (EDU) Division of Research on Learning in Formal and Informal (DRL).
@@ Appendix A Training Details
Our models are trained largely following the procedure in Misra & Mahowald (2024). We have a total of eleven model-types, of which nine are discussed in the main text (with two following a more strict definition of datives discussed in App.˜D), and for each type, we conducted three training runs with different seeds to account for variation due to initialization. In total, this amounts to 33 LM training runs. See Table˜3 for a summary of training details—hyperparameters, compute hardware, etc.
@@ C.2 Global Regression Models
In Table˜4, the interaction coefficients (e.g., balanced : Length Difference) can be thought of as telling us how much the Length or Animacy difference differs in the given condition, as opposed to the default baseline.
@@
Intercept 0.459-0.459***
balanced 0.0460.046***
no-datives 0.0780.078***
no-2postverbal 0.0870.087***
swapped-datives 0.1210.121***
short-first 0.1270.127***
random-first 0.1390.139***
long-first 0.2490.249***
long-first-headfinal 0.2480.248***
Length Difference 0.189-0.189***
Animacy Difference 0.0910.091***
balanced : Length Difference 0.1050.105***
no-datives : Length Difference 0.1350.135***
no-2postverbal : Length Difference 0.1400.140***
swapped-datives : Length Difference 0.2270.227***
short-first : Length Difference 0.1310.131***
random-first : Length Difference 0.1870.187***
long-first : Length Difference 0.2250.225***
long-first-headfinal : Length Difference 0.2660.266***
balanced : Animacy Difference 0.117-0.117***
no-datives : Animacy Difference 0.104-0.104***
no-2postverbal : Animacy Difference 0.109-0.109***
swapped-datives : Animacy Difference 0.195-0.195***
short-first : Animacy Difference 0.070-0.070***
random-first : Animacy Difference 0.066-0.066***
long-first : Animacy Difference 0.034-0.034*
long-first-headfinal : Animacy Difference 0.126-0.126
@@
long-first-headfinal 0.1090.109***
@@
short-first : Length difference 0.0550.055***
long-first : Length Difference 0.0390.039***
long-first-headfinal : Length Difference 0.0800.080***
@@
long-first : Animacy Difference 0.0320.032*
long-first-headfinal : Animacy Difference 0.0600.060***
@@ Appendix D Strict-dative default and balanced Results
Table 6: Regression coefficients from mixed effect model, fit separately for strict default and balanced predicting DO preference, with *** for $p<.001$.
@@
Both the regression coefficients in Table˜6 and the length and animacy effects in Fig.˜7 for the strict models are similar to their loose counterparts. We therefore focus on just the loose notion in the main paper, so that the ablated models are more aggressively removing direct evidence.
@@ Appendix E Verb-specific Random Slopes for Animacy
Random slopes by verb lemma for default, balanced, swapped-datives, and no-datives are shown in Fig.˜8. We see that animacy effects for certain verbs are heavily affected by the manipulations in balanced and swapped-datives, indicating that the animacy preferences of these verbs are more tied to direct evidence. However, no-datives recovers similar, albeit a bit weaker, animacy effects as in default. This suggests that usages of dative verbs in the indirect evidence can inform an LM’s dative animacy preferences.
Figure 8: Verb-specific random slope-differences (differences from the global random slope) for the animacy term in our linear mixed-effects model analysis. Positive value indicates relatively greater DO preference for animate recipients, while negative ones indicate lower DO preference for inanimate recipients.
@@ Appendix F Detecting Datives and 2postverbals
Using spaCy, we first detect all utterances that could contain a DO or a PO dative. For DOs, we search for utterances containing a verb which has a dative and a dobj dependency or two dobj dependencies. For POs, we look for utterances containing a verb with a dobj dependency, and with either a to prepositional phrase under a dative or prep dependency or a for prepositional phrase under a dative dependency. From these entries, we perform a sanity check to ensure that the noun phrases come after the verb, and the recipient and theme have appropriate parts of speech.
These entries are then loose-datives. From these entries, we check if the verb lemma is classified as a dative or benefactive verb in Levin (1993). If so, we check if it is in the permissible form (e.g. if the verb is classified as benefactive only, the entry cannot be a to-PO, or if the verb is classified as DO only, the entry cannot be a PO). If a loose-dative satisfies the above, it is also a strict-dative.
@@ Appendix G Artificial Pollution
The counts of dative exposure by model are listed in Table˜7.
@@ Appendix I Measuring the Short-Firstness of English
For each sentence, we compute the inversion number corresponding to a head that has $n\geq 2$ dependents as the number of out-of-order pairs (w.r.t. to short-first or long-first). The inversion number of a head is precisely equal to the number of adjacent swaps to sort its dependents. As such, a head with $n$ dependents has an inversion number at most $\binom{n}{2}$.
@@
For entries where constituent rearrangements were possible (i.e. there exists one head with two or more dependents), it takes 23.8% of the maximal number of swaps to reach short-first, compared to 53.2% for long-first. This means that, on average, it takes more than twice the number of adjacent-swap operations to sort by long-first compared to short-first.